# SuPreME: A Supervised Pre-training Framework for Multimodal ECG Representation Learning

### v2

@@ Abstract
Cardiovascular diseases are a leading cause of death and disability worldwide. Electrocardiogram (ECG) is critical for diagnosing and monitoring cardiac health, but obtaining large-scale annotated ECG datasets is labor-intensive and time-consuming. Recent ECG Self-Supervised Learning (eSSL) methods mitigate this by learning features without extensive labels but fail to capture fine-grained clinical semantics and require extensive task-specific fine-tuning. To address these challenges, we propose SuPreME, a Supervised Pre-training framework for Multimodal ECG representation learning. SuPreME is pre-trained using structured diagnostic labels derived from ECG report entities through a one-time offline extraction with Large Language Models (LLMs), which help denoise, standardize cardiac concepts, and improve clinical representation learning. By fusing ECG signals with textual cardiac queries instead of fixed labels, SuPreME enables zero-shot classification of unseen conditions without further fine-tuning. We evaluate SuPreME on six downstream datasets covering 106 cardiac conditions, achieving superior zero-shot AUC performance of 77.20%, surpassing state-of-the-art eSSLs by 4.98%. Results demonstrate SuPreME’s effectiveness in leveraging structured, clinically relevant knowledge for high-quality ECG representations.
@@
Mingsheng Cai1,3, Jiuming Jiang1,3, Wenhao Huang2, Che Liu3thanks: Correspondence: che.liu21@imperial.ac.uk, Rossella Arcucci3 1The University of Edinburgh, 2Shenzhen Yinwang Intelligent Technology Co., Ltd, 3Imperial College London {m.cai, jiuming.jiang}@ed.ac.uk1, huangwenhao@yinwang.com2, {mingsheng.cai23, jiuming.jiang23, che.liu21, r.arcucci}@imperial.ac.uk3
@@ 2 Related Work
ECG Supervised Learning. ECG supervised learning (eSL) methods, using CNNs or Transformers in Figure 2(a--b), achieve high accuracy in cardiovascular disease diagnosis. CNNs excel at capturing spatial and temporal patterns in 1D ECG signals or 2D ECG images (Tesfai et al., 2022; Degirmenci et al., 2022; Mashrur et al., 2019; Huang et al., 2022), while Transformers use attention mechanisms to model global dependencies (Natarajan et al., 2020; Jiang et al., 2021; He et al., 2023). Despite their strengths, eSLs rely heavily on large-scale datasets with expert-verified annotations, making them costly and impractical for pre-training tasks (Strodthoff et al., 2020). This dependence limits their scalability and generalizability, particularly when addressing diverse datasets or unseen cardiac conditions.
ECG Self-supervised Learning. To overcome the annotation bottleneck, ECG self-supervised learning (eSSL) methods have been introduced, enabling representation learning from unannotated ECG signals in Figure 2(c--d). Contrastive learning frameworks, such as CLOCS and ASTCL (Kiyasseh et al., 2021; Wang et al., 2023), explore temporal and spatial invariance in ECG data (Eldele et al., 2021; Chen et al., 2020, 2021). Generative eSSL techniques reconstruct masked segments to capture signal-level features (Zhang et al., 2022; Sawano et al., 2022; Na et al., 2024; Jin et al., ). Despite their successes, eSSLs fail to incorporate clinical semantics from associated medical reports and require fine-tuning for downstream tasks (Liu et al., 2023d, c; He et al., 2022), limiting their utility in zero-shot scenarios.
@@ 3.2 Multimodal ECG Supervised Learning
[…]
@@ 4.1 Configuration and Settings
Clinical Entity Extraction. Following Section 3.1, we extract and normalize clinical entities from MIMIC-IV-ECG using Llama3.1-70B-Instruct [^2] Used offline for one-time inference only; not required during deployment. A large LLM ensures high-quality labels. (Appendix A.10.1) with structured prompts to ensure high-quality annotations. Entities are deduplicated via MedCPT embeddings (cosine similarity $>0.8$) and mapped to UMLS/SNOMED CT (average cosine similarity $>0.75$) [^3] Verified by cardiologists with 10+ years of experience; Appendix A.10.3.. Experiments are run on 8 NVIDIA A100-SMX4-80GB GPUs using vLLM Kwon et al. (2023). Statistics of extracted MIMIC-IV-ECG entities are in Appendix A.6.1.
Supervised ECG Pre-training. We use a 1D ViT-tiny encoder (patch size = 125, i.e., 0.25s) and a frozen MedCPT text encoder. Training employs AdamW (LR=$1\times 10^{-3}$, weight decay=$1\times 10^{-8}$) with cosine annealing ($T_{0}$=5000, $T_{\text{mult}}$=1, min LR=$1\times 10^{-8}$), for up to 50 epochs with early stopping (patience=10, best AUC at 16). Batch size is set to 256 on 4 NVIDIA A100-PCIE-40GB GPUs [^4] Compact pre-trained checkpoint (ViT-tiny + frozen MedCPT) runs on single GPU with $\geq$24GB memory, making it deployable in clinical or low-resource settings (Appendix A.10.2)..
@@ 4.2 Evaluation with Mainstream eSSLs
To investigate SuPreME’s sensitivity to pre-training scale, we evaluate its zero-shot performance under varying data proportions (Figure 5). SuPreME consistently improves with more data and maintains a clear advantage over non-multimodal eSSLs. Remarkably, with only 20% of pre-training data, SuPreME outperforms all non-multimodal eSSLs using 1% or 10% labeled data for linear probing, and matches the zero-shot performance of the multimodal baseline MERL trained on 100% of the pre-training data. Moreover, SuPreME’s zero-shot performance with just 20% of data also exceeds MERL’s linear probing result with 1% labels, highlighting its superior generalization and efficiency under limited supervision. Notably, SuPreME achieves these results with significantly fewer computational resources and shorter training time [^5] SuPreME: 4$\times$×A100-40GB GPUs for 16 epochs ($\sim$90 minutes); MERL: 8$\times$×A100-40GB GPUs for 50 epochs ($\geq$1 day). (Appendix A.10.2).
@@ 4.3 Evaluation of SuPreME Architecture
Figure 7 further analyzes SuPreME’s performance on individual cardiac conditions in PTB-XL-Subclass (Complete results in Appendix A.9). SuPreME consistently achieves high AUCs (many $>90$), especially for nuanced conditions like LAFB/LPFB, CRBBB, CLBBB, and RVH, where both query semantics and signal patterns must be integrated. In contrast, ResNet + CFN underperforms in complex arrhythmias (e.g., ISCA, IRBBB), reinforcing our insight that strong multimodal fusion requires compatible encoders.
@@ A.1.2 Downstream Dataset
CPSC-2018. The CPSC-2018888CPSC-2018 is available at http://2018.icbeb.org/Challenge.html. dataset originates from the China Physiological Signal Challenge (CPSC) 2018, including 6,877 records from 9,458 patients, with durations ranging from 6 to 60 seconds Liu et al. (2018). The standard 12-lead ECG data is sampled at a rate of 500 Hz, collected from 11 hospitals and categorized into 9 different labels: 1 normal type and 8 abnormal types.
@@ A.3.3 Case Study of Deduplication and Mapping
NORM normal Normal (1.000), of normal (0.995), Normal result (0.979), Normal interest (0.953), percent of normal (0.953), percent of normal (0.942)
@@ A.5.2 Simplified Clinical Knowledge-Enhanced Prompt Engineering
“A conduction abnormality characterized by delayed depolarization of the left ventricle, typically resulting in a widened QRS complex ($>$120 ms), often associated with underlying structural heart disease or ischemia.”
@@ A.5.3 Evaluation Metrics
AUROC is the area under the ROC curve, with values ranging from 0 to 1, reflecting the overall classification ability of the model. AUROC = 0.5 indicates that the model’s classification ability is equivalent to random guessing, while AUROC $>$ 0.5 and values closer to 1 indicate that the model is able to classify with greater accuracy.
@@ A.10.1 Offline Use of Large-sized LLM for Clinical NER
While we employ LLaMA3.3-70B-Instruct for clinical entity extraction, this step is performed offline only once during dataset construction and is not part of the SuPreME model’s pre-training & inference pipeline. The motivation for using a larger model is to ensure high annotation quality for the pre-training dataset. Once entities are extracted and mapped, they form a standardized query list used throughout training and evaluation. Therefore, clinical deployments do not require access to large LLMs, and the SuPreME model itself remains lightweight during inference.
@@ A.10.2 Computation Cost and Practical Deployment
Data Processing (Offline NER). To obtain high-quality supervision labels, we extract and normalize diagnostic entities from MIMIC-IV-ECG reports using LLaMA3.3–70B-Instruct with structured prompts. This step is performed only once as described above before pre-training SuPreME model. The output is a cleaned, deduplicated dataset of standardized diagnostic labels, which serves as supervision for training. The annotation process takes approximately 6 hours on 8 NVIDIA A100-SMX4-80GB GPUs, and the tested minimum reproducing resources are 4 NVIDIA A100-PCIE-40GB GPUs without parallelized LLM inference. The resulting standardized dataset is reused across training and downstream evaluation.
@@ A.10.3 Similarity Threshold Determination
A.10.4 Comparing SuPreME with MERL
@@
A.10.5 On the Effectiveness of CFN with Different Backbones
@@
A.10.6 Domain Scope and Generalization Potential

### v3

@@ Abstract
Cardiovascular diseases are a leading cause of death and disability worldwide. Electrocardiogram (ECG) is critical for diagnosing and monitoring cardiac health, but obtaining large-scale annotated ECG datasets is labor-intensive and time-consuming. Recent ECG Self-Supervised Learning (eSSL) methods mitigate this by learning features without extensive labels but fail to capture fine-grained clinical semantics and require extensive task-specific fine-tuning. To address these challenges, we propose SuPreME, a Supervised Pre-training framework for Multimodal ECG representation learning. SuPreME is pre-trained using structured diagnostic labels derived from ECG report entities through a one-time offline extraction with Large Language Models (LLMs), which help denoise, standardize cardiac concepts, and improve clinical representation learning. By fusing ECG signals with textual cardiac queries instead of fixed labels, SuPreME enables zero-shot classification of unseen conditions without further fine-tuning. We evaluate SuPreME on six downstream datasets covering 106 cardiac conditions, achieving superior zero-shot AUC performance of 77.20%, surpassing state-of-the-art eSSLs by 4.98% [^1] All code and data are available at https://github.com/mingscai/SuPreME.. Results demonstrate SuPreME’s effectiveness in leveraging structured, clinically relevant knowledge for high-quality ECG representations.
@@
Mingsheng Cai1,2, Jiuming Jiang1,2, Wenhao Huang3, Che Liu2thanks: Correspondence: che.liu21@imperial.ac.uk, Rossella Arcucci2 1The University of Edinburgh, 2Imperial College London, 3Shenzhen Yinwang Intelligent Technology Co., Ltd {mingsheng.cai, jiuming.jiang}@ed.ac.uk1, {mingsheng.cai23, jiuming.jiang23, che.liu21, r.arcucci}@imperial.ac.uk2, huangwenhao@yinwang.com3
@@ 2 Related Work
ECG Supervised Learning. ECG supervised learning (eSL) methods, using CNNs or Transformers in Figure 2(a-b), achieve high accuracy in cardiovascular disease diagnosis. CNNs excel at capturing spatial and temporal patterns in 1D ECG signals or 2D ECG images (Tesfai et al., 2022; Degirmenci et al., 2022; Mashrur et al., 2019; Huang et al., 2022), while Transformers use attention mechanisms to model global dependencies (Natarajan et al., 2020; Jiang et al., 2021; He et al., 2023). Despite their strengths, eSLs rely heavily on large-scale datasets with expert-verified annotations, making them costly and impractical for pre-training tasks (Strodthoff et al., 2020). This dependence limits their scalability and generalizability, particularly when addressing diverse datasets or unseen cardiac conditions.
ECG Self-supervised Learning. To overcome the annotation bottleneck, ECG self-supervised learning (eSSL) methods have been introduced, enabling representation learning from unannotated ECG signals in Figure 2(c-d). Contrastive learning frameworks, such as CLOCS and ASTCL (Kiyasseh et al., 2021; Wang et al., 2023), explore temporal and spatial invariance in ECG data (Eldele et al., 2021; Chen et al., 2020, 2021). Generative eSSL techniques reconstruct masked segments to capture signal-level features (Zhang et al., 2022; Sawano et al., 2022; Na et al., 2024; Jin et al., ). Despite their successes, eSSLs fail to incorporate clinical semantics from associated medical reports and require fine-tuning for downstream tasks (Liu et al., 2023d, c; He et al., 2022), limiting their utility in zero-shot scenarios.
@@ 3.2 Multimodal ECG Supervised Learning
Let $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{M}\}$ denote the global query list. Each query $q_{i}$ is encoded into a dense vector using the query encoder from MedCPT, which applies a Transformer (Trm) to the input sequence $\text{CLS}]\ q_{i}\ [\text{SEP}]$. The final-layer $\text{CLS}]$ token embedding is used as the query representation:
@@ 4.1 Configuration and Settings
Clinical Entity Extraction. Following Section 3.1, we extract and normalize clinical entities from MIMIC-IV-ECG using Llama3.1-70B-Instruct [^3] Used offline for one-time inference only; not required during deployment. A large LLM ensures high-quality labels. (Appendix A.10.1) with structured prompts to ensure high-quality annotations. Entities are deduplicated via MedCPT embeddings (cosine similarity >0.8>0.8) and mapped to UMLS/SNOMED CT (average cosine similarity >0.75>0.75) [^4] Verified by cardiologists with 10+ years of experience; Appendix A.10.3,  A.10.4.. Experiments are run on 8 NVIDIA A100-SMX4-80GB GPUs using vLLM Kwon et al. (2023). Statistics of extracted MIMIC-IV-ECG entities are in Appendix A.6.1.
[…]
@@ 4.2 Evaluation with Mainstream eSSLs
[…]
@@
Uncertainty Filtering Zero-shot AUC
Not Filtered 63.96 $\pm$ 0.47
Filtered (Ours) 77.20 $\pm$ 0.21
Table 5: Uncertain entity filtering.
@@ 4.3 Evaluation of SuPreME Architecture
Figure 7 further analyzes SuPreME’s performance on individual cardiac conditions in PTB-XL-Subclass (Complete results in Appendix A.9). SuPreME consistently achieves high AUCs (many >90>90), especially for nuanced conditions like LAFB/LPFB, CRBBB, CLBBB, and RVH, where both query semantics and signal patterns must be integrated. In contrast, ResNet + CFN underperforms in complex arrhythmias (e.g., ISCA, IRBBB), reinforcing our insight that strong multimodal fusion requires compatible encoders.
@@ 4.4 Ablation Analysis
Uncertain Entity Filtering. A cardiac-specific vocabulary with synonym consolidation ensures reliable alignment (Table 12). Removing the uncertainty filtering step lowers zero-shot AUC from 77.20% to 63.96% (Table 5), confirming its importance.
@@ A.1.2 Downstream Dataset
CPSC-2018. The CPSC-2018999CPSC-2018 is available at http://2018.icbeb.org/Challenge.html. dataset originates from the China Physiological Signal Challenge (CPSC) 2018, including 6,877 records from 9,458 patients, with durations ranging from 6 to 60 seconds Liu et al. (2018). The standard 12-lead ECG data is sampled at a rate of 500 Hz, collected from 11 hospitals and categorized into 9 different labels: 1 normal type and 8 abnormal types.
@@ A.3.3 Case Study of Deduplication and Mapping
NORM normal Normal (1.000), of normal (0.995), Normal result (0.979), Normal interest (0.953), percent of normal (0.942)
@@ A.5.2 Simplified Clinical Knowledge-Enhanced Prompt Engineering
“A conduction abnormality characterized by delayed depolarization of the left ventricle, typically resulting in a widened QRS complex (>>120 ms), often associated with underlying structural heart disease or ischemia.”
@@ A.5.3 Evaluation Metrics
AUROC is the area under the ROC curve, with values ranging from 0 to 1, reflecting the overall classification ability of the model. AUROC = 0.5 indicates that the model’s classification ability is equivalent to random guessing, while AUROC >> 0.5 and values closer to 1 indicate that the model is able to classify with greater accuracy.
@@ A.10.1 Offline Use of Large-sized LLM for Clinical NER
While we employ LLaMA3.1-70B-Instruct for clinical entity extraction, this step is performed offline only once during dataset construction and is not part of the SuPreME model’s pre-training & inference pipeline. The motivation for using a larger model is to ensure high annotation quality for the pre-training dataset. Once entities are extracted and mapped, they form a standardized query list used throughout training and evaluation. Therefore, clinical deployments do not require access to large LLMs, and the SuPreME model itself remains lightweight during inference.
@@ A.10.2 Computation Cost and Practical Deployment
Data Processing (Offline NER). To obtain high-quality supervision labels, we extract and normalize diagnostic entities from MIMIC-IV-ECG reports using LLaMA3.1–70B-Instruct with structured prompts. This step is performed only once as described above before pre-training SuPreME model. The output is a cleaned, deduplicated dataset of standardized diagnostic labels, which serves as supervision for training. The annotation process takes approximately 6 hours on 8 NVIDIA A100-SMX4-80GB GPUs, and the tested minimum reproducing resources are 4 NVIDIA A100-PCIE-40GB GPUs without parallelized LLM inference. The resulting standardized dataset is reused across training and downstream evaluation.
@@ A.10.3 Similarity Threshold Determination
A.10.4 LLM-based NER Evaluation
Regarding the performance of NER outputs from LLaMA, we conducted a manual evaluation on 200 randomly sampled clinical reports, annotated by physicians with over ten years of experience. Using a token-level BIO tagging scheme (Ramshaw and Marcus, 1999; Moryossef et al., 2023), we found strong agreement between LLaMA and expert annotations, with an F1 score of 95.71% and a Cohen’s kappa of 0.93 (Cohen, 1960; Ruan et al., 2024). These results confirm the reliability of LLaMA for clinical NER and support the validity of our original findings.
Metrics Value
Precision 96.30%
Recall 95.13%
F1 95.71%
Cohen’s kappa 0.93
Table 21: Physician-annotated evaluation of LLaMA NER performance.
A.10.5 Comparing SuPreME with MERL
@@
To ensure completeness, we also compare SuPreME with MERL-ResNet in zero-shot settings. Across six datasets, SuPreME achieves higher AUCs on four tasks and yields an overall average gain of +1.95 absolute AUC (equivalent to +2.6% relative improvement).
A.10.6 On the Effectiveness of CFN with Different Backbones
@@
A.10.7 Domain Scope and Generalization Potential